### v1

# Exploring How Audio Effects Alter Emotion with Foundation Models

Abstract
Audio effects (FX) such as reverberation, distortion, modulation, and dynamic range processing play a pivotal role in shaping emotional responses during music listening. While prior studies have examined links between low-level audio features and affective perception, the systematic impact of audio FX on emotion remains underexplored. This work investigates how foundation models—large-scale neural architectures pretrained on multimodal data—can be leveraged to analyze these effects. Such models encode rich associations between musical structure, timbre, and affective meaning, offering a powerful framework for probing the emotional consequences of sound design techniques. By applying various probing methods to embeddings from deep learning models, we examine the complex, nonlinear relationships between audio FX and estimated emotion, uncovering patterns tied to specific effects and evaluating the robustness of foundation audio models. Our findings aim to advance understanding of the perceptual impact of audio production practices, with implications for music cognition, performance, and affective computing.
Index Terms—  Foundation Models, Affect Computing, Audio FX, Model Auditing
1 Introduction
The relationship between sound and human emotion has been extensively studied across multiple domains, particularly within affective computing and music cognition. Prior research has demonstrated how low-level audio features—such as timbre, tempo, pitch, and rhythm—correlate with emotional responses in listeners [1, 2, 3, 4, 5]. Parallel investigations have examined how environmental auditory conditions, including soundscape complexity and the number of simultaneous sound sources, influence affective perception [6]. However, there remains a notable gap in understanding how audio FX—such as reverberation, distortion, modulation, and dynamic range processing—systematically alter emotional responses in music listening contexts.
Unlike fundamental audio features, audio FX are intentional sound design tools commonly applied in music production and performance to modulate aesthetic and emotional impact. Yet, their role in shaping affective perception has been underexplored. For instance, [7] showed that longer reverberation times in recorded music increase ratings of “sublimity,” including emotional qualities like awe and nostalgia. Similarly, distortion has been linked to heightened arousal and affective intensity, potentially evoking excitement or even aggression [2]. Higher rates of vibrato are linked to elevated arousal levels [8, 9]. Additionally, an increase in the number of perceived sound sources correlates with increased arousal but may reduce pleasantness [6].
Despite these isolated findings, there is no comprehensive study that systematically investigates how diverse audio FX in music—individually or in combination—impact emotional perception. Given that audio FX are central to contemporary music production, a focused analysis of their affective consequences is both timely and necessary.
At the same time, advances in foundation models—large-scale neural architectures pre-trained on multimodal datasets, have demonstrated an ability to capture high-level semantic and perceptual relationships across domains, including with audio FX [10, 11]. In the music domain, foundation models such as MusicLM [12], MERT [13], Qwen [14], and CLAP [15] have shown promising abilities in generating emotionally expressive audio, learning disentangled latent representations, and enabling cross-modal retrieval tasks. These models implicitly encode rich associations between musical structure, timbre, and affective meaning, making them ideal tools to probe novel hypotheses about music-emotion relationships.
In this work, we leverage foundation models to investigate how specific audio FX influence emotional perception in music. We analyze performance changes, shifts in predicted emotional labels and dimensions, and alterations in model embeddings across three state-of-the-art foundation models and diverse datasets with both categorical and dimensional annotations, using six common audio FX at different intensity levels, including fixed real-world. By combining controlled manipulations with these realistic production scenarios, we identify correlations between audio FX and estimated emotions and uncover properties of how foundation models represent these affective cues.
Table 1: Audio FX $\Delta$ (Level10-Level1) across MSE Arousal, MSE Valence, and F1 score. Bold values indicate the largest absolute difference among audio FXs for a specific model and dataset, while underlined values indicate the largest absolute difference among models for a specific audio FX and dataset.
Audio FX MSE Arousal MSE Valence F1 Score Categorical
DEAM witheFlow DEAM witheFlow witheFlow EMOPIA
Qwen MERT CLAP Qwen MERT CLAP Qwen MERT CLAP Qwen MERT CLAP Qwen MERT CLAP Qwen MERT CLAP
Chorus +0.003 +0.008 -0.001 +0.060 +0.110 +0.029 +0.007 +0.004 -0.001 +0.120 +0.120 +0.070 -0.160 -0.190 -0.090 -0.310 -0.110 -0.070
Distortion +0.012 +0.023 +0.028 +0.160 +0.122 -0.027 +0.009 +0.017 +0.027 +0.240 +0.210 +0.110 -0.488 -0.433 -0.149 -0.370 -0.390 -0.350
Phaser +0.006 +0.005 +0.002 +0.171 +0.129 -0.040 +0.006 +0.006 +0.001 +0.075 +0.168 +0.030 -0.266 -0.230 -0.041 -0.180 -0.100 -0.020
EQ +0.007 +0.010 +0.014 +0.051 +0.060 -0.033 +0.005 +0.007 +0.011 +0.050 +0.140 -0.050 -0.103 -0.186 +0.041 +0.040 -0.060 +0.030
Delay +0.003 +0.004 -0.001 +0.016 +0.113 +0.008 +0.005 +0.005 -0.002 +0.060 +0.060 +0.040 -0.091 -0.147 +0.000 -0.200 -0.090 +0.100
Reverb +0.021 +0.012 +0.010 -0.012 +0.141 +0.106 +0.020 +0.015 +0.006 +0.030 +0.060 +0.055 -0.022 -0.150 -0.116 -0.160 -0.230 -0.170
2 Material
In this study, we employed a diverse set of datasets and models, spanning both deep and shallow architectures, along with several commonly used audio FXs with varying parameter settings. The code, analytical details, and complete experimental results are available in our GitHub repository  [^1] https://github.com/stelioskt/audioFX.
We utilized three datasets capturing both categorical and dimensional emotional annotations: EMOPIA [16], DEAM [17], and witheFlow [5]. The EMOPIA dataset is a multimodal (audio and MIDI) resource focused on perceived emotion in pop piano music, containing 1,087 clips from 387 songs. Each clip is annotated with one of four categorical emotions—Excitement, Anger, Sadness, or Calmness—supporting both clip-level classification and song-level emotion analysis. The DEAM dataset comprises 1,802 excerpts and full songs annotated with valence and arousal as continuous, time-varying values per second, as well as aggregated ratings, along with metadata on duration, genre, and folksonomy tags, enabling regression-based modeling of emotion dynamics. The part of the witheFlow dataset we used includes 235 solo-instrument recordings, annotated by 20 listeners with continuous valence-arousal values per second following the Circumplex Model of Affect [18], as well as multi-label categorical tags based on the Geneva Emotional Music Scale (GEMS-9) [19], making it suitable for both dimensional and categorical emotion modeling in music.
(a) CLAP-Chorus
(b) CLAP-Delay
(c) CLAP-Distortion
(d) Qwen-Chorus
(e) Qwen-Delay
(f) Qwen-Distortion
(g) MERT-Chorus
(h) MERT-Delay
(i) MERT-Distortion
Fig. 1: Radar plots of emotion predictions for CLAP, Qwen, and MERT across three audio effects for the EMOPIA dataset. Each level in the plots depicts the distribution of emotions, normalised based on the greatest value in a given plot.
To investigate and interpret musical emotion, we employed three state-of-the-art foundation models, each offering unique capabilities. First, MERT-v1-330M [13] is a 330M-parameter model (24 layers with 1024 hidden units) trained on 160,000 hours of music audio using masked language modeling with RVQ-VAE and CQT teacher signals. MERT operates on 24 kHz audio at 75 frames per second and is particularly suited to capturing fine-grained musical and acoustic emotion cues. Second, CLAP [20] is a dual-encoder model with roughly 630M parameters, combining a Transformer-based (HTS-AT) audio encoder with a RoBERTa-based text encoder. CLAP was trained on 128K–630K audio–text pairs via contrastive learning, using log-Mel spectrogram inputs at 48 kHz to learn a shared audio–text representation space with strong zero-shot generalization. Finally, Qwen2-Audio-7B [21] is a 7B-parameter model featuring a Whisper-style audio encoder and Qwen-style decoder. Qwen2-Audio-7B was trained on over 300K hours of speech, music, and environmental audio to support broad audio-language tasks, including automatic speech recognition, classification, question answering, and conversational interaction.
For the head classifiers of our foundation models, we employed shallow interpretable models tailored to each task and dataset, training them while keeping the foundation model frozen. For regression on DEAM and witheFlow valence-arousal annotations, we used an XGBRegressor [22] to predict continuous emotional dimensions, evaluating performance with MAE, MSE, and $R^{2}$. For single-label classification on EMOPIA, we applied an XGBClassifier, assessing weighted accuracy, precision, recall, and F1-score. For multi-label classification on witheFlow categorical tags, we implemented a OneVsRest strategy with XGBClassifier, measuring F1-micro and F1-macro scores with a threshold of 0.5. This configuration allowed us to effectively adapt gradient-boosted tree models to both dimensional and categorical emotion prediction tasks.
Fig. 2: UMAP visualization of foundation model embeddings for the EMOPIA dataset, showing trajectories generated for each input identity after applying audio FX with the intensity ranging from 1 to 10.
To apply controlled alterations to our input and observe the behavior of our chosen foundation model, we used six audio FX from the pedalboard library, each scaled in intensity from 1 to 10. Reverb simulates acoustic space by adjusting room_size, creating the perception of different environments from small rooms to large halls, which can influence temporal and spectral characteristics of the signal. Delay introduces echoes through delay_seconds and feedback, affecting rhythmic perception and reinforcing certain frequencies. Distortion adds harmonic saturation via drive_db, enhancing harmonic content and introducing nonlinearities that challenge the model’s robustness. EQ shapes the frequency spectrum with low_cutoff and high_cutoff filters, emphasizing or attenuating spectral bands and altering timbral balance. Chorus thickens the sound by duplicating and slightly detuning the signal with rate, depth, and feedback, producing subtle modulation that can affect perceived pitch and texture. Phaser applies sweeping frequency modulation via rate, depth, and feedback, generating moving notches that alter harmonic relationships over time.
3 Experiments
Using the above material, we conducted four experiments to investigate the relationships between audio effects and estimated emotions. First, we analyzed performance changes to assess how audio manipulations impact model accuracy. Second, we examined shifts in emotion predictions to identify changes in predicted emotional labels or dimensions. Third, we explored embedding space alterations to understand how audio effects influence the internal representations learned by the models. Finally, we investigated how real-world audio FX settings affect emotion estimation.
3.1 Impact on Model Performance
Inspired by [10], that studies the impact of audio effects on foundation models, and [23], that examines performance changes in speech emotion recognition models induced by codec processing, we aimed to explore how incorporating emotional information affects performance in emotion recognition tasks. To investigate this, we trained interpretable machine learning classifiers using embeddings extracted from foundation models across all datasets. This approach allowed us to develop models for multi-class classification, multi-label classification, and regression tasks.
After training our interpretable classifiers, we applied various audio effects at different intensity levels and analyzed their impact on model performance. All foundation models were consistently used as backbones across tasks and datasets. We hypothesized that any observed performance degradation would indicate a potential correlation between the audio FX and the estimated emotion.
Overall, we observed a general decrease in performance, indicating that audio effects may influence emotion recognition. As summarized in Table 1, phaser and distortion in particular evoked the largest performance declines across multiple intensity levels. All three models exhibited similar trends, with the decrease becoming more pronounced as the intensity of the audio effects increased.
3.2 Shifts in Emotion Predictions
Previously trained pipelines were used to run inference on audio samples that had been modified with six common effects applied at varying intensities. Transparent classifiers built on different foundation model backbones generated emotion predictions from these altered inputs (see Section 2). The resulting predictions were compared with those from the original samples, and the observed changes were visualized to reveal how the effects shifted predicted emotions. This approach enabled the examination of potential correlations between specific audio manipulations and estimated emotional responses.
In Fig. 1, we examine how emotional predictions in the EMOPIA dataset are affected by different audio effects. Notably, high levels of distortion consistently increase Anger predictions while decreasing Calmness across all models [24]. In contrast, delay and chorus effects introduce greater variability, suggesting that these manipulations create ambiguity in the models’ emotional interpretations [25]. Interestingly, increasing the chorus effect tends to boost Calmness predictions in the CLAP and MERT models, whereas increasing the delay effect raises Anger predictions in the CLAP and Qwen models.
3.3 Embedding Space Alterations
In this section, we present an experiment investigating how the embeddings of a foundation model—representations that encode emotional information—are altered as common audio effects are gradually applied. Inspired by the visualization approach of Deng et al. [10], we identified the most important features from our shallow emotion classification model. Using the top 25 features, we performed UMAP [26] dimensionality reduction and visualized the foundation model’s embeddings for inputs modified at different effect intensities. This approach allowed us to observe how audio effects reshape the structure of the embeddings and to identify which models are most sensitive to these changes.
To ensure stability and comparability, embeddings were first standardized per (dataset, model) using zero-mean/unit-variance scaling, then cleaned via (i) variance thresholding (removing features with variance $1\times 10^{-6}$) and (ii) iterative correlation pruning (dropping features with $|r|>0.95$). Feature selection was then performed: for regression-style datasets (DEAM, witheFlow) we used ElasticNetCV with $l1\_ratio\in\{0.5,0.8\}$, $\alpha$ values log-spaced from $10^{-3}$ to $10^{2}$ (50 values), 5-fold cross-validation, $\max\_iter=60000$, and $tol=2\times 10^{-3}$, while for classification-style datasets (EMOPIA, witheFlow) we used logistic regression with an elastic net penalty (l1_ratio=0.5l1\_ratio=0.5), the SAGA solver, $C=0.5$, max_iter=6000max\_iter=6000, and $tol=2\times 10^{-3}$. For visualization, we applied UMAP using a cosine metric and spectral initialization; in this experiment, these shallow models were chosen for their linearity and interpretability, with $K=25$, n_neighbors=30n\_neighbors=30, and min_dist=0.5min\_dist=0.5. To avoid crowding and label imbalance, we sampled 20 tracks per effect for regression datasets, 5 tracks per label for classification datasets, and 3 tracks per label for the multi-label dataset (witheFlow), using a fixed random seed ($42$), covering the effects reverb, delay, distortion, EQ, chorus, and phaser.
In Fig. 2, we show how embeddings from each foundation model respond to audio effects on the EMOPIA dataset. Our analysis reveals systematic differences in how foundation models encode the impact of audio effects. CLAP embeddings exhibit large, structured displacements as FX levels increase—particularly for delay, chorus, and distortion—indicating strong sensitivity to timbral manipulations. Qwen also shows noticeable shifts, though with less consistent trajectories, whereas MERT remains relatively stable across all effect levels, suggesting robustness to such manipulations, likely due to its training on music-specific tasks. Overall, trajectory length and variance in embedding space can serve as metrics for quantifying the influence of audio FX on emotion, with the magnitude of this impact depending strongly on the choice of foundation model.
3.4 Real World Scenario
We tested real-world effect chains designed to mimic iconic sounds from Pink Floyd, U2 (emphasizing reverb and delay to create atmospheric textures), and Rage Against the Machine (heavy distortion with moderate chorus)  [^2] https://www.kitrae.net/music/David_Gilmour_Effects_And_Gear.html [^3] https://www.guitarchalk.com/guitar-amp-settings-guide/#t-1649265629386. The goal of this experiment was to examine whether, in real-world scenarios, carefully crafted audio FX chains can produce stronger shifts in emotion predictions due to their intentional artistic design.
As shown in Fig. 3, moving from isolated effects to complete FX chains highlights how cumulative production choices shift embeddings in latent space. MERT and Qwen show similar, structured trajectories reflecting these cumulative effects. The distortion-heavy chain of Rage Against the Machine produces nearly unidirectional shifts, while the reverb- and delay-heavy chains of U2 and Pink Floyd show more complex, diffuse paths. CLAP exhibits shorter, scattered movements, indicating reduced sensitivity to combined effects. Overall, real-world FX chains induce larger, more coherent shifts, suggesting that trajectory patterns—unidirectional versus multidirectional—can serve as proxies for how effect categories shape emotional perception.
Fig. 3: UMAP visualization of foundation model embeddings for the witheFlow dataset, showing trajectories generated for each input after applying audio FX with the intensity ranging from 1 to 10.
4 Conclusions and Future Work
Our study demonstrates that audio effects substantially alter estimated emotion in music. Distortion and phaser, in particular, strongly increase Anger while reducing Calmness, whereas chorus and delay introduce more variability in predictions. Analysis of embedding-space trajectories shows that the magnitude and structure of latent-space shifts reflect the emotional impact of FX, with real-world effect chains producing the most pronounced and coherent changes, likely because they are intentionally designed by artists to evoke strong emotional responses. Among the models, MERT exhibited relative robustness, while CLAP and Qwen were more sensitive to manipulations.
For future work, we plan to investigate the simultaneous application of multiple audio effects, extend the study to additional datasets and foundation models, explore alternative probing techniques to gain deeper insight into how complex audio transformations influence emotional perception in music, and include comparisons to classical feature-based systems as well as analyses of human ground truth after effects.